OpenStudy (anonymous):

simplify 49^(-1/2)power

OpenStudy (anonymous):

well this means simply\[\sqrt{49^{-1}}\]

OpenStudy (anonymous):

A 1/2 power is a square root, and a negative power is in the denominator, so 49^(-1/2) is \[\frac{ 1 }{ \sqrt{49} }\]

OpenStudy (anonymous):

in order to remove the - from the equation you must flip 49 \[\sqrt{\frac{ 1 }{ 49 }}\] which simplifies to \[\pm \frac{ 1 }{ 7 }\]

OpenStudy (anonymous):

ok thank u